# PLeaS — Merging Models with Permutations and Least Squares

### v1

@@ Acknowledgement
This work is supported by Microsoft Grant for Customer Experience Innovation and the National Science Foundation under grant no. 2019844, 2112471, and 2229876. JH is supported by the NSF Graduate Research Fellowship Program.
@@ Appendix A Experimental and Implementation details
For finding permutation symmetries, we use the official implementation of Git Re-Basin at this url. We also rely on the implementation of ZipIt! for the comparisons in Section 5.
For solving the least squares objective for PLeaS , we use SGD with a batch size of 32, a learning rate of $10^{-3}$. We sample equally from both datasets in each batch for experiments involving data. We run our algorithm for 100 steps, and find that it converges quickly. For PLeaS-Act , we similarly compute the activations on 100 batches of data for matching and finding the optimal permutations. We also reset batch norm parameters using 100 batches of data from the actual domains for all methods.
@@ A.1 Compute time and cost
A.2 Computing the layerwise merging ratio
Note that $k_{i}$ can be different for each layer. Given a configuration $K=\{\frac{k_{i}}{d_{i}}:i\in[L]\}$, we can model the FLOPs/memory of the merged model as a quadratic function of $k_{i}$, which we denote as $\text{Footprint}(K)$. For a given relative memory/FLOPs budget $B$, we want to find $K$ s.t. $\text{Footprint}(K)\leq B$ to maximize the accuracy of a model merged with the configuration $K$. We scale everything so that $B=1$ corresponds to the footprint of a single model. This problem is NP-Hard. We propose a relaxation of the problem in order to get an approximate solution. First, we measure the performance of a set of models merged with “leave one out" configurations of $K$, where for each layer $i$, we construct $K_{i}^{0}=\{k_{j}:k_{j}=d_{i}\ \text{if}\ j=i,\ 0\ \text{otherwise}\}$ and $K_{i}^{1}=\{k_{j}:k_{j}=0\ \text{if}\ j=i,\ d_{i}\ \text{otherwise}\}$. $K_{i}^{0}$ corresponds to merging only layer $i$, keeping all other layers unmerged, and $K_{i}^{1}$ corresponds to merging every other layer while keeping $i$ unmerged. We also compute the accuracies of the fully merged model (denoted by $K^{0}$) and the ensemble (denoted by $K^{1}$). Then, we approximate the accuracy of any given $K$ with a linear function as
$\text{Acc}(K)=\sum_{i=1}^{L}\frac{k_{i}}{d_{i}}((2-B)(\text{Acc}(K^{1})-\text{Acc}(K_{i}^{0}))-(1-B)(\text{Acc}(K_{i}^{0})-\text{Acc}(K^{0})))$
@@ B.1 What to optimize for Least Squares?
In Eq. 1, we propose to solve a least squares problem involving the permuted average activations from each layer of the component models. In Table 3, we demonstrate that this choice is not only natural, but also performs better than other alternatives. It is also interesting to note that the second row in the table corresponds to a permuted version of RegMean[14]. This formulation performs better than RegMean, indicating that using permutations is necessary to align features for networks which were differently initialized. Further, row 3 is similar to the objective proposed by [11], but we show that PLeaS outperforms this objective as well.
B.2 Merging models trained on CIFAR and CINIC
In this section, we present results on merging ResNet-18 models trained on CIFAR-10 and CINIC-Im datasets. CINIC-Im is a subset of the CINIC-10 [6] dataset which does not contain any images from CIFAR-10. We follow the same experimental protocol outlined in Section 5 to train two independent models on these datasets from scratch. We then merge the models and report the results for merging models with 1×1$\times$1 × size of ResNet-18 in Table 4. We make a few surprising observations in this case. We find that randomly permuting the weights of one model and averaging these weights achieves a non-trivial performance, and this is improved by using Git Re-Basin to find the optimal permutation. Using PLeaS further improves this performance.
Table 4: Accuracy while merging models trained on CIFAR and CINIC-Im: We find that random permutations achieve a non-trivial performance, and PLeaS can improve the overall accuracy.
Method CIFAR CINIC-Im
Ensemble 94.5 79.1
Random Permutation 55.1 62.7
Git Re-Basin 67.2 68.3
PLeaS 77.4 77.6
B.3 Merging ResNet-18
(a) Shared Label Space
(b) Different Label Spaces
Figure 5: Memory-Performance trade-off for ResNet-18: We merge pairs of models fine-tuned on different datasets, and compute the average performance across all four datasets for two settings - datasets with a Shared label space (Fig. 5(a)), and datasets with Different label spaces (Fig. 5(b))
In Fig. 5, we plot the performance of PLeaS against baselines for merging pairs of ResNet-18 under the same settings as Section 5. We find that PLeaS outperforms the baselines at lower FLOPs, and the difference reduces for larger FLOP budgets.
B.4 Reducing the accuracy barrier on ImageNet
@@
In this section, we show the performance of PLeaS while merging ResNet-50 models trained independently on ImageNet. The accuracy of a single model on this task is 77.5%. As seen from Fig 6(a), current methods including ZipIt![28] and Git Re-Basin[1] struggle on merging models for this task, with the accuracy of the merged model being significantly lower than the accuracy of a single model. This has been referred to as the accuracy barrier on ImageNet in prior work. PLeaS makes some progress towards lowering this barrier, and improves over Git Re-Basin by over 9% at 1.0×1.0$\times$1.0 × FLOPs budget. For context, this accuracy is at par with that obtained by merging WideResNet-50 models with a width multiplier of 2 using Git Re-Basin. More promisingly, the flexibility afforded by partially permuting and merging models gives another avenue to lower the accuracy barrier, with a model of size 1.4×1.4$\times$1.4 × having an accuracy barrier of 2% with PLeaS . However, further work is needed to reduce this accuracy barrier. In Fig. 6(b), we compare using synthetic data from [3] for all purposes of activation computation while merging ImageNet trained models. We find that using PLeaS-free with synthetic data can come close to using actual data, being within 1% in terms of accuracy at 1.2×1.2$\times$1.2 × model size.
B.5 Detailed Results
Each of our evaluation was run across three random restarts. These random restarts shuffle the data used for computing activations and merging the models. They also affect the initialization of the merged model. Each pair evaluation was also run twice, swapping the order of pre-trained models used for either of the datasets of the pair. We hence have 6 runs for each dataset pair. In Tables 5 and 6, we provide the results for each dataset pair, reporting the average and standard deviation across the 6 runs.
Table 5: Detailed Results on DomainNet
@@ B.5 Detailed Results
Table 6: Detailed Results on Different Label Spaces
@@
PLeaS 1.2 Imagenet 71.0 ±$\pm 0.2$ 88.1 ±$\pm 0.2$ 68.5 ±$\pm 1.0$ 70.1 ±$\pm 0.5$ 79.2 ±$\pm 0.6$ 70.4 ±$\pm 0.3$ 73.7 ±$\pm 1.2$ 73.4 ±$\pm 0.7$ 73.5 ±$\pm 1.8$ 89.1 ±$\pm 0.5$ 77.5 ±$\pm 0.6$ 89.9 ±$\pm 0.3$
PLeaS 1.2 Original 72.5 ±$\pm 0.2$ 88.5 ±$\pm 0.2$ 70.9 ±$\pm 0.5$ 72.5 ±$\pm 0.5$ 80.5 ±$\pm 0.4$ 72.5 ±$\pm 0.3$ 75.2 ±$\pm 0.8$ 76.4 ±$\pm 0.4$ 77.2 ±$\pm 0.9$ 90.7 ±$\pm 0.6$ 79.9 ±$\pm 0.3$ 91.5 ±$\pm 0.9$
PLeaS 1.0 Imagenet 68.6 ±$\pm 0.5$ 79.4 ±$\pm 1.0$ 66.2 ±$\pm 0.8$ 58.9 ±$\pm 0.6$ 76.5 ±$\pm 0.6$ 65.8 ±$\pm 1.0$ 69.1 ±$\pm 1.3$ 63.9 ±$\pm 0.7$ 71.7 ±$\pm 1.1$ 83.7 ±$\pm 0.7$ 70.9 ±$\pm 0.4$ 84.6 ±$\pm 0.7$
PLeaS 1.0 Original 70.6 ±$\pm 0.3$ 79.2 ±$\pm 0.4$ 69.6 ±$\pm 0.6$ 62.2 ±$\pm 0.4$ 79.4 ±$\pm 0.6$ 69.8 ±$\pm 0.4$ 72.4 ±$\pm 0.7$ 69.1 ±$\pm 0.5$ 75.2 ±$\pm 0.8$ 87.0 ±$\pm 0.4$ 76.1 ±$\pm 0.4$ 89.4 ±$\pm 0.5$
PLeaS 1.8 Imagenet 73.4 ±$\pm 0.6$ 91.7 ±$\pm 0.5$ 71.2 ±$\pm 0.7$ 75.8 ±$\pm 0.5$ 80.0 ±$\pm 0.6$ 72.2 ±$\pm 0.2$ 75.5 ±$\pm 0.8$ 77.2 ±$\pm 0.5$ 76.1 ±$\pm 0.7$ 92.0 ±$\pm 0.7$ 80.9 ±$\pm 0.5$ 91.9 ±$\pm 0.6$
PLeaS 1.8 Original 74.8 ±$\pm 0.2$ 91.6 ±$\pm 0.2$ 73.9 ±$\pm 0.3$ 77.5 ±$\pm 0.6$ 81.7 ±$\pm 0.4$ 74.8 ±$\pm 0.4$ 76.5 ±$\pm 0.5$ 80.2 ±$\pm 0.3$ 78.8 ±$\pm 0.4$ 92.3 ±$\pm 0.5$ 82.8 ±$\pm 0.5$ 92.9 ±$\pm 0.5$
PLeaS 1.55 Imagenet 72.5 ±$\pm 0.2$ 89.9 ±$\pm 0.1$ 70.1 ±$\pm 1.0$ 74.1 ±$\pm 0.8$ 79.4 ±$\pm 0.6$ 71.6 ±$\pm 0.3$ 74.9 ±$\pm 0.6$ 75.8 ±$\pm 0.7$ 75.8 ±$\pm 1.1$ 91.0 ±$\pm 0.7$ 79.8 ±$\pm 0.7$ 91.4 ±$\pm 0.5$
PLeaS 1.55 Original 73.9 ±$\pm 0.4$ 90.0 ±$\pm 0.2$ 72.8 ±$\pm 0.5$ 76.0 ±$\pm 0.6$ 81.4 ±$\pm 0.3$ 74.0 ±$\pm 0.3$ 76.1 ±$\pm 0.5$ 78.7 ±$\pm 0.5$ 77.9 ±$\pm 0.7$ 91.9 ±$\pm 0.4$ 81.7 ±$\pm 0.5$ 92.3 ±$\pm 0.7$
@@
Permutation 1.2 Original 67.9 ±$\pm 0.7$ 87.5 ±$\pm 0.7$ 66.8 ±$\pm 0.3$ 69.6 ±$\pm 0.4$ 78.2 ±$\pm 0.5$ 68.3 ±$\pm 0.3$ 72.0 ±$\pm 0.8$ 72.3 ±$\pm 0.5$ 73.9 ±$\pm 0.9$ 88.9 ±$\pm 0.3$ 77.2 ±$\pm 0.7$ 89.7 ±$\pm 0.7$
Permutation 1.0 Original 60.6 ±$\pm 0.4$ 78.1 ±$\pm 0.5$ 57.0 ±$\pm 0.7$ 59.1 ±$\pm 0.5$ 72.3 ±$\pm 0.6$ 59.9 ±$\pm 0.3$ 59.8 ±$\pm 0.8$ 60.4 ±$\pm 0.8$ 66.5 ±$\pm 0.6$ 82.1 ±$\pm 0.8$ 67.0 ±$\pm 0.7$ 82.5 ±$\pm 0.5$
Permutation 1.8 Original 74.0 ±$\pm 0.6$ 91.0 ±$\pm 0.7$ 73.0 ±$\pm 0.4$ 76.3 ±$\pm 0.5$ 81.2 ±$\pm 0.3$ 73.8 ±$\pm 0.2$ 76.0 ±$\pm 0.5$ 78.5 ±$\pm 0.5$ 77.7 ±$\pm 0.6$ 92.1 ±$\pm 0.6$ 82.2 ±$\pm 0.4$ 92.5 ±$\pm 1.0$
Permutation 1.55 Original 71.4 ±$\pm 0.6$ 90.4 ±$\pm 0.5$ 71.1 ±$\pm 0.4$ 73.9 ±$\pm 0.5$ 80.1 ±$\pm 0.5$ 71.9 ±$\pm 0.3$ 75.0 ±$\pm 0.7$ 76.2 ±$\pm 0.4$ 76.5 ±$\pm 0.9$ 91.0 ±$\pm 0.5$ 80.4 ±$\pm 0.3$ 91.9 ±$\pm 0.5$
@@
B.6 Using task specific heads

### v2

@@ Acknowledgement
This work is supported by Microsoft Grant for Customer Experience Innovation and the National Science Foundation under grant no. 2019844, 2112471, and 2229876. JH is supported by the NSF Graduate Research Fellowship Program. PWK is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Innovation under the AI Visiting Professorship Programme (award number AIVP-2024-001).
@@ Appendix A Experimental and Implementation details
For finding permutation symmetries, we use the official implementation of Git Re-Basin at this url. We also rely on the implementation of ZipIt! and MuDSC for the comparisons in Sec. 5.
For solving the least squares objective for PLeaS, we use SGD with a batch size of 32, a learning rate of $10^{-3}$. We sample equally from both datasets in each batch for experiments involving data. We run our algorithm for 100 steps, and find that it converges quickly. For the first step of PLeaS, we similarly compute the activations on 100 batches of data for matching and finding the optimal permutations. We also reset batch norm parameters using 100 batches of data from the actual domains for all methods.
@@
As an example evaluation for Domainnet, we have 8 models, two each on Clipart, Infograph, Painting and Real domains. We merge these pair-wise. We hence have 12 (6 domain pairs and two models per pair) merged models. For each merged model, we compute the accuracy on its component domains. Hence, for each domain, we have 6 performances (3 domain pairs and two models per pair). We report the average of these 6 numbers in Tab. 1.
ViT
To obtain models for merging with CIFAR-50+50 for Sec. 5.5.1, we follow the protocol from [32]. In particular, we use ImageNet pretrained ViT models, and train two such models on disjoint 50 class subsets of CIFAR-100. These are trained with CLIP language embeddings for the final layer. This process is repeated thrice to get different sets of classes and pairs of models. After merging, the performance of the merged model is measured on CIFAR-100 as well as the two subsets of classes that the moders were trained on. The average of these results is reported in Tab. 3. For the other datasets, we start off with ViT-B/16 model from the OpenCLIP project[11] which is pretrained on the LAION-400M dataset. We fix the text encoder, and fine-tune the image encoder on various datasets. This fine-tuned model achieves an accuracy of (72%, 91%, 68%, 24%) on CUB, Pets, Dogs and NABirds datasets resp. We then merge the model pairwise and report the results.
@@ A.1 Compute time and cost
A.2 Computing the permutation matrices
We use the algorithms of Git Re-Basin to compute the permutation matrices $P_{i}$. For activation matching, we collect the representations for both models over a batch of data, and measure the alignment between two neurons as the squared distance between their representations (closer $\rightarrow$ better alignment). Then for each layer, we find the bipartite matching (i.e. permutation) between the two models that minimizes the total distance using a standard algorithm (scipy.optimize.linear_sum_assignment). For weight matching, the alignment between two neurons is the squared distance between its input and output vectors. This time, the permutations at adjacent layers interact, so we perform an alternating minimization, solving for the permutations one layer at a time until we reach a fixed point.
A.3 Computing the layerwise merging ratio
Note that $k_{i}$ can be different for each layer. Given a configuration $K=\{\frac{k_{i}}{d_{i}}:i\in[L]\}$, we can model the FLOPs/memory of the merged model as a quadratic function of $k_{i}$, which we denote as $\text{Footprint}(K)$. For a given relative memory/FLOPs budget $B$, we want to find $K$ s.t. $\text{Footprint}(K)\leq B$ to maximize the accuracy of a model merged with the configuration $K$. We scale everything so that $B=1$ corresponds to the footprint of a single model. This problem is NP-Hard. We propose a relaxation of the problem in order to get an approximate solution. First, we measure the performance of a set of models merged with “leave one out” configurations of $K$, where for each layer $i$, we construct $K_{i}^{0}=\{k_{j}:k_{j}=d_{i}\ \text{if}\ j=i,\ 0\ \text{otherwise}\}$ and $K_{i}^{1}=\{k_{j}:k_{j}=0\ \text{if}\ j=i,\ d_{i}\ \text{otherwise}\}$. $K_{i}^{0}$ corresponds to merging only layer $i$, keeping all other layers unmerged, and $K_{i}^{1}$ corresponds to merging every other layer while keeping $i$ unmerged. We also compute the accuracies of the fully merged model (denoted by $K^{0}$) and the ensemble (denoted by $K^{1}$). Then, we approximate the accuracy of any given $K$ with a linear function as
$\text{Acc}(K)=\sum_{i=1}^{L}\frac{k_{i}}{d_{i}}\bigg{(}(2-B)(\text{Acc}(K^{1})-\text{Acc}(K_{i}^{0}))$
$-(1-B)(\text{Acc}(K_{i}^{0})-\text{Acc}(K^{0}))\bigg{)}$
@@ A.3 Computing the layerwise merging ratio
A.3.1 Empirical Results
(a) PLeaS
(b) Permutations
Figure 5: Comparing our strategy for layer-wise merging with a linear baseline: We merge models using PLeaS  and permutations using the strategy described in Sec. A.3 and a linear strategy where $\frac{k}{d}$ is held constant.
In Figs. 5(a) and 5(b), we compare the QP method with a baseline strategy which assigns the number of units in each layer to be a constant ratio. We find that our strategy outperforms this baseline for both PLeaS  and Permutations.
@@ B.1 What to optimize for Least Squares?
In Eq. 1, we propose to solve a least squares problem involving the permuted average activations from each layer of the component models. In Tab. 4, we demonstrate that this choice is not only natural, but also performs better than other alternatives. It is also interesting to note that the second row in the table corresponds to a permuted version of RegMean[14]. This formulation performs better than RegMean, indicating that using permutations is necessary to align features for networks which were differently initialized. Further, row 3 is similar to the objective proposed by [10], but we show that PLeaS outperforms this objective as well.
B.2 Reducing the accuracy barrier on ImageNet
@@
In this section, we show the performance of PLeaS while merging ResNet-50 models trained independently on ImageNet. The accuracy of a single model on this task is 77.5%. As seen from Fig. 6(a), current methods including ZipIt! [28] and Git Re-Basin [1] struggle on merging models for this task, with the accuracy of the merged model being significantly lower than the accuracy of a single model. This has been referred to as the accuracy barrier on ImageNet in prior work. PLeaS makes some progress towards lowering this barrier, and improves over Git Re-Basin by over 9% at 1.0×1.0$\times$1.0 × FLOPs budget. For context, this accuracy is at par with that obtained by merging WideResNet-50 models with a width multiplier of 2 using Git Re-Basin. More promisingly, the flexibility afforded by partially permuting and merging models gives another avenue to lower the accuracy barrier, with a model of size 1.4×1.4$\times$1.4 × having an accuracy barrier of 2% with PLeaS. However, further work is needed to reduce this accuracy barrier. In Fig. 6(b), we compare using synthetic data from [6] for all purposes of activation computation while merging ImageNet trained models. We find that using PLeaS-free with synthetic data can come close to using actual data, being within 1% in terms of accuracy at 1.2×1.2$\times$1.2 × model size.
B.3 Detailed Results
Each of our evaluation was run across three random restarts. These random restarts shuffle the data used for computing activations and merging the models. They also affect the initialization of the merged model. Each pair evaluation was also run twice, swapping the order of pre-trained models used for either of the datasets of the pair. We hence have 6 runs for each dataset pair. In Tabs. 5 and 6, we provide the results for each dataset pair, reporting the average and standard deviation across the 6 runs.
Table 5: Detailed Results on DomainNet We report the results for ResNet-50 here
@@ B.3 Detailed Results
Table 6: Detailed Results on Different Label Spaces We report the results for ResNet-50 here
@@
PLeaS-Weight 1.2 Imagenet 71.0 ±$\pm 0.2$ 88.1 ±$\pm 0.2$ 68.5 ±$\pm 1.0$ 70.1 ±$\pm 0.5$ 79.2 ±$\pm 0.6$ 70.4 ±$\pm 0.3$ 73.7 ±$\pm 1.2$ 73.4 ±$\pm 0.7$ 73.5 ±$\pm 1.8$ 89.1 ±$\pm 0.5$ 77.5 ±$\pm 0.6$ 89.9 ±$\pm 0.3$
PLeaS-Weight 1.2 Original 72.5 ±$\pm 0.2$ 88.5 ±$\pm 0.2$ 70.9 ±$\pm 0.5$ 72.5 ±$\pm 0.5$ 80.5 ±$\pm 0.4$ 72.5 ±$\pm 0.3$ 75.2 ±$\pm 0.8$ 76.4 ±$\pm 0.4$ 77.2 ±$\pm 0.9$ 90.7 ±$\pm 0.6$ 79.9 ±$\pm 0.3$ 91.5 ±$\pm 0.9$
PLeaS-Weight 1.0 Imagenet 68.6 ±$\pm 0.5$ 79.4 ±$\pm 1.0$ 66.2 ±$\pm 0.8$ 58.9 ±$\pm 0.6$ 76.5 ±$\pm 0.6$ 65.8 ±$\pm 1.0$ 69.1 ±$\pm 1.3$ 63.9 ±$\pm 0.7$ 71.7 ±$\pm 1.1$ 83.7 ±$\pm 0.7$ 70.9 ±$\pm 0.4$ 84.6 ±$\pm 0.7$
PLeaS-Weight 1.0 Original 70.6 ±$\pm 0.3$ 79.2 ±$\pm 0.4$ 69.6 ±$\pm 0.6$ 62.2 ±$\pm 0.4$ 79.4 ±$\pm 0.6$ 69.8 ±$\pm 0.4$ 72.4 ±$\pm 0.7$ 69.1 ±$\pm 0.5$ 75.2 ±$\pm 0.8$ 87.0 ±$\pm 0.4$ 76.1 ±$\pm 0.4$ 89.4 ±$\pm 0.5$
PLeaS-Weight 1.8 Imagenet 73.4 ±$\pm 0.6$ 91.7 ±$\pm 0.5$ 71.2 ±$\pm 0.7$ 75.8 ±$\pm 0.5$ 80.0 ±$\pm 0.6$ 72.2 ±$\pm 0.2$ 75.5 ±$\pm 0.8$ 77.2 ±$\pm 0.5$ 76.1 ±$\pm 0.7$ 92.0 ±$\pm 0.7$ 80.9 ±$\pm 0.5$ 91.9 ±$\pm 0.6$
PLeaS-Weight 1.8 Original 74.8 ±$\pm 0.2$ 91.6 ±$\pm 0.2$ 73.9 ±$\pm 0.3$ 77.5 ±$\pm 0.6$ 81.7 ±$\pm 0.4$ 74.8 ±$\pm 0.4$ 76.5 ±$\pm 0.5$ 80.2 ±$\pm 0.3$ 78.8 ±$\pm 0.4$ 92.3 ±$\pm 0.5$ 82.8 ±$\pm 0.5$ 92.9 ±$\pm 0.5$
PLeaS-Weight 1.55 Imagenet 72.5 ±$\pm 0.2$ 89.9 ±$\pm 0.3$ 70.1 ±$\pm 1.0$ 74.1 ±$\pm 0.8$ 79.4 ±$\pm 0.6$ 71.6 ±$\pm 0.3$ 74.9 ±$\pm 0.6$ 75.8 ±$\pm 0.7$ 75.8 ±$\pm 1.1$ 91.0 ±$\pm 0.7$ 79.8 ±$\pm 0.7$ 91.4 ±$\pm 0.5$
PLeaS-Weight 1.55 Original 73.9 ±$\pm 0.4$ 90.0 ±$\pm 0.2$ 72.8 ±$\pm 0.5$ 76.0 ±$\pm 0.6$ 81.4 ±$\pm 0.3$ 74.0 ±$\pm 0.3$ 76.1 ±$\pm 0.5$ 78.7 ±$\pm 0.5$ 77.9 ±$\pm 0.7$ 91.9 ±$\pm 0.4$ 81.7 ±$\pm 0.5$ 92.3 ±$\pm 0.7$
@@
Permutation-Weight 1.2 Original 67.9 ±$\pm 0.7$ 87.5 ±$\pm 0.7$ 66.8 ±$\pm 0.3$ 69.6 ±$\pm 0.4$ 78.2 ±$\pm 0.5$ 68.3 ±$\pm 0.3$ 72.0 ±$\pm 0.8$ 72.3 ±$\pm 0.5$ 73.9 ±$\pm 0.9$ 88.9 ±$\pm 0.3$ 77.2 ±$\pm 0.7$ 89.7 ±$\pm 0.7$
Permutation-Weight 1.0 Original 60.6 ±$\pm 0.4$ 78.1 ±$\pm 0.5$ 57.0 ±$\pm 0.7$ 59.1 ±$\pm 0.5$ 72.3 ±$\pm 0.6$ 59.9 ±$\pm 0.3$ 59.8 ±$\pm 0.8$ 60.4 ±$\pm 0.8$ 66.5 ±$\pm 0.6$ 82.1 ±$\pm 0.8$ 67.0 ±$\pm 0.7$ 82.5 ±$\pm 0.5$
Permutation-Weight 1.8 Original 74.0 ±$\pm 0.6$ 91.0 ±$\pm 0.7$ 73.0 ±$\pm 0.4$ 76.3 ±$\pm 0.5$ 81.2 ±$\pm 0.3$ 73.8 ±$\pm 0.2$ 76.0 ±$\pm 0.5$ 78.5 ±$\pm 0.5$ 77.7 ±$\pm 0.6$ 92.1 ±$\pm 0.6$ 82.2 ±$\pm 0.4$ 92.5 ±$\pm 1.0$
Permutation-Weight 1.55 Original 71.4 ±$\pm 0.6$ 90.4 ±$\pm 0.5$ 71.1 ±$\pm 0.4$ 73.9 ±$\pm 0.5$ 80.1 ±$\pm 0.5$ 71.9 ±$\pm 0.3$ 75.0 ±$\pm 0.7$ 76.2 ±$\pm 0.4$ 76.5 ±$\pm 0.9$ 91.0 ±$\pm 0.5$ 80.4 ±$\pm 0.3$ 91.9 ±$\pm 0.5$
@@
B.4 Using task specific heads